OpenStudy (anonymous):

The equation of a line is shown below: y = - x - 7 What is the equation of a line which is parallel to this line and passes through (1, 79)? A. y = -x + 80 B. y = x - 80 C. y = 7x - 80 D. y = -7x + 80

OpenStudy (anonymous):

@bluewaffles

OpenStudy (anonymous):

@Lyrae @RolyPoly @shaadykhamlichi

OpenStudy (lyrae):

The lines are parallel if they have the same slope.

OpenStudy (lyrae):

y = mx + b m is the slope. Which line in the answers have the same slope (m) as y = -x - 7 ?

OpenStudy (anonymous):

B right ? @Lyrae

OpenStudy (lyrae):

Does B y = 1x - 80 have the same m as y = (-1)x + 80 ?

OpenStudy (anonymous):

i think so :(

OpenStudy (lyrae):

y = 1x - 80 y = (-1)x + 80 ^ ^ | | m m m = 1 m = -1

OpenStudy (anonymous):

darn

OpenStudy (lyrae):

The (minus) sign is important!

OpenStudy (lyrae):

Any idea which one it might be now?

OpenStudy (anonymous):

A :))))

OpenStudy (lyrae):

Yep! :)

OpenStudy (anonymous):

thank you a lot